# Cosmological Zoom-In Simulations of Milky Way Host Size Dark Matter Halos with a Blue-Tilted Primordial Power Spectrum

### v1

@@ III.2.1 Cosmological parameters
    Root-mean-square matter fluctuation averaged over a sphere of radius $8h^{-1}~{}\rm{Mpc}$, $\sigma_{8}=0.807$.
@@ III.3 Halo finder and halo property post-processor
In the main text of the paper, we adopt HBT+HERONS (Han et al., 2017) (Forouhar Moreno et al in preparation)  [^5] This is an updated version of HBT+ that improves the tracking of subhalos. It can be found at https://github.com/SWIFTSIM/HBT-HERONS as our halo finder, which is a tracking finder. We also explore the result with a phase space finder, VELOCIraptor (Elahi et al., 2019), and compare their results in Appendix A. In this paper, we adopt 20 as the minimum number of particles to be identified as a dark matter halo for both halo finders. We choose the position of the most bound particle as the halo center.
We use SOAP (Spherical Overdensity and Aperture Processor) [^6] https://github.com/SWIFTSIM/SOAP to calculate the spherical overdensity properties, based on the results from HBT+HERONS.
@@ IV.1 Terminology
    where $R$ is the distance to the halo center, $M(<R)$ the enclosed mass that is bound to the halo, within radius $R$. The $R-V_{\rm cir}(R)$ graph is called the galaxy rotation curve, and $V_{\rm max}$ is simply the maximum circular velocity of the rotation curve. In the simulation, the subhalo would lose its mass due to the tidal forces from the host halo. However, $V_{\rm max}$ is robust and representative of a dark matter subhalo because it reflects the subhalo’s central, gravitationally bound part, which is less affected by the stripping effect and how the boundary of the subhalo is defined.
@@ IV.2 Comparison between the PL and the BT models
Figure 4: The comparison between the dark matter mass projection maps with the PL and BT (deep/soft) PPS for m12i at z=0. The color scales are the same across the three panels. All the images are 2D plane projections of cubic volumes with a side length of $400~{}\rm{kpc}$, centered at the main halo’s most bound particle identified by HBT+HERONS.
Table 2 shows some selected properties for the main halos in all the simulations. We learn that the main halo of BT_soft is larger than that of the PL model in mass, radius, concentration, and $V_{\rm max}$. However, the BT_deep model behaves similarly to the PL model, consistent with the pivot mass $M_{p}$ based on Equation 5.
@@
In the rest of this subsection, we would quantify these differences in terms of the subhalo mass function, the subhalo $V_{\rm max}$ function (velocity refers to $V_{\rm max}$ of the subhalos), the subhalo radial distribution (subhalo distance function), the cumulative substructure mass fraction, the subhalo $R_{\rm max}-V_{\rm max}$ relationship, and the main halo radial density profile.
@@ IV.2.1 Main Halo Radial Density Profile
The upper panels of Figure 5 show a good match between the simulation results and the NFW fitting lines. Thus we can use the concentration parameters given by the NFW profile to describe the radial density profile. The concentration parameters of BT_deep and PL are pretty close (see Table 2), consistent with their similar radial density profiles for the main halos. Meanwhile, the BT_soft model gets a larger concentration parameter than others, consistent with its higher radial density profiles. The unresolved regions are based on the criterion described by subsubsection III.3.1.
@@ IV.2.2 Subhalo Mass Function
Here, $m$ indicates the maximum relative ratio the BT model can touch compared to the PL model, and $\mu_{c}$ is the middle point between $\mu$ starting to drop and $\mu$ value finishing the dropping. According to the definition of Equation 8, $\mu_{c}$ is also where the inverse S shape function drops to half. Since the deep and soft models perform pretty closely in the subhalo mass function, we can use one set of parameters to fit: $m=2,\mu_{c}=10^{-3}$. After getting this fitting function from the bottom ratio panel, we draw it on the upper panel of Figure 6, showing a good fit with the simulation results.
@@ IV.2.4 Subhalo Radial Distribution
As suggested by Springel et al. (2008), within the range of $10^{5}~{}{\rm M}_{\odot}\sim 10^{10}~{}{\rm M}_{\odot}$, massive subhalos tend to exist in the radially outer region while different mass bins follow the same line. Both our BT and PL simulation results also show this trend. Although the BT subhalo number density is higher, the radial range where subhalos exist within that mass range does not change.
@@ V CONCLUSIONS
JWST’s high redshift observations revealed that galaxies could form earlier than those in the standard model of cosmology Labbé et al. (2023); Boylan-Kolchin (2023); Lovell et al. (2023). The strong lensing observations prefer more substructures Macciò and Miranda (2006); Xu et al. (2009, 2015). Furthermore, the recent observations Müller et al. (2024); Homma et al. (2024) and theories Kim et al. (2018) suggested there could be a “too many satellites” problem in the nearby universe. These motivate us to consider alternative cosmological models with an enhanced Primordial Power Spectrum at small scales ($>1~{}{\rm Mpc}^{-1}$), namely the blue-tilted (BT) model.
In this study, we have conducted cosmological zoom-in simulations of MW host-size halos with the blue-tilted primordial power spectrum. We have considered two blue-tilted models from Ref. Hirano and Yoshida (2024): BT_soft with the enhancement at $\gtrsim 1~{}{\rm Mpc}^{-1}$ and BT_deep with the enhancement at $\gtrsim 4~{}{\rm Mpc}^{-1}$. They are all within the accepted parameter space according to the JWST observation Hirano and Yoshida (2024) and the dwarf galaxy central density observation Dekker and Kravtsov (2024).
@@ Acknowledgments
TKC is supported by the ‘Improvement on Competitiveness in Hiring New Faculties’ Funding Scheme from the Chinese University of Hong Kong (4937210, 4937211, 4937212). TKC and JW are partially supported by the Direct Grant project (4053662) from the Chinese University of Hong Kong. VJFM acknowledges support by NWO through the Dark Universe Science Collaboration (OCENW.XL21.XL21.025).
@@ Appendix A Results from another halo finder: VELOCIraptor
Apart from HBT+HERONS, we also identify the halos and subhalos with VELOCIraptor (VR) (Elahi et al., 2019), a code based on the friends-of-friends (FOF) algorithm that identifies peaks in 6D phase-space to find subhalos (both the configuration space and phase space). We compare the HBT+HERONS and VR results on the subhalo mass function, subhalo $V_{\rm max}$ function, and subhalo radial distribution in subsection A.1.
Currently, SOAP (see subsection III.3) only allows the calculation of spherical overdensity properties for field halos. Besides, HBT+HERONS (without SOAP post process) could only calculate limited kinds of SO properties. Hence, we used VR (without SOAP post process) to calculate the virial masses for all the halos (also including subhalos) and show the $M_{\rm vir}-V_{\rm max}$ relationship in subsection A.2. Usually, the spherical overdensity properties will include all of the particles within the SO radius, regardless of whether they are bound to the main halo. However, for subhalos, VR only considers the bound particles [^8] See Mass and radius properties in https://velociraptor-stf.readthedocs.io/en/latest/output.html. Otherwise, the subhalos near the center of the main halo will have unphysically large masses.
Our essential parameters for VR are listed in footnote 11, based on the sample configuration for 6D dark matter-only simulation  [^9] https://velociraptor-stf.readthedocs.io/en/latest/_downloads/97ba16973bcff9d4afe940c2e13d12d2/sample_dmcosmological_6dfof_subhalo.cfg. We set the most bound particle’s position as the reference frame center to be consistent with HBT+HERONS for comparison in subsection A.1.
@@
A.1 Comparison between halo finders: HBT+HERONS and VELOCIraptor
Here, we choose m12i to analyze the difference between the two halo finders. VR stands for VELOCIraptor in both the text and figures. In this subsection, we also use SOAP to process VR and HBT+HERONS catalogs in a consistent manner.
@@ A.1 Comparison between halo finders: HBT+HERONS and VELOCIraptor
Figure 11: The cumulative scaled subhalo mass functions for m12i, with two halo finders HBT+HERONS (colored solid) and VR (colored dashed), respectively. The upper panel shows the subhalo mass function for the PL (orange), BT_deep (green), and BT_soft (red) models, with the subhalo mass $M_{\rm sub}$ scaled by $M_{200c}$ of the main halo. The bottom panel shows the ratios of the VR to HBT+HERONS number. The shaded areas are the same as that in Figure 6. The shaded areas overlap because we set the minimum size of the subhalo to be the same in both halo finders.
The subhalo mass function comparison between VR and HBT+HERONS is shown in Figure 11. These two halo finders agree with each other for $10^{-5}<\mu<10^{-3},10^{7}~{}{\rm M}_{\odot}<M_{\rm sub}<10^{9}~{}{\rm M}_{\odot}$. However, HBT+HERONS is higher and closer to the COCO fitting line than VR at the high mass end, implying that HBT+HERONS is better at finding more massive subhalos. But at the low mass end, VR appears to have more subhalos than HBT+HERONS.
Figure 12: The cumulative scaled subhalo $V_{\rm max}$ functions for m12i, with two halo finders HBT+HERONS (colored solid) and VR (colored dashed), respectively. The upper panel shows the subhalo $V_{\rm max}$ function for the PL (orange), BT_deep (green), and BT_soft (red) models, with the subhalo $V_{\rm max}$ scaled by $V_{200c}$ of the main halo. The bottom panel shows the ratios of the VR to HBT+HERONS number. The shaded areas are the same as those in Figure 7.
The subhalo $V_{\rm max}$ function comparison between VR and HBT+HERONS is shown in Figure 12. These two halo finders give similar results except at the low $V_{\rm max}$ end near the resolution limit. When approaching the limit $V_{\rm max}$ value, VR deviates from the power-law behavior later than HBT+HERONS.
Figure 13: The subhalo radial number density profiles at four different mass bins for m12i, using different halo finders HBT+HERONS (colored solid) and VR (colored dashed). We consider the PL (orange), BT_deep (green), and BT_soft (red) models in each subfigure (also in each mass bin): the upper panel is the same definition as that in Figure 8; the bottom panel shows the ratios of the VR results to the HBT+HERONS results.
The subhalo radial distribution comparison between VR and HBT+HERONS is shown in Figure 13. In mass bins larger than $10^{7}~{}{\rm M}_{\odot}$, VR and HBT+HERONS have nearly the same results. But, at the lower mass end, VR has much fewer subhalos in the inner region than HBT+HERONS and the Aquarius result.
To summarize, VR and HBT+HERONS behave similarly for the subhalo mass and $V_{\rm max}$ functions. However, for the radial properties (the subhalo radial distribution function), VR gaves a lower halo count near the host halo center for the low halo mass bin.

### v2

@@ III.2.1 Cosmological parameters
    Root-mean-square matter fluctuation averaged over a sphere of radius $8h^{-1}~{}\rm{cMpc}$, $\sigma_{8}=0.807$.
@@ III.3 Halo finder and halo property post-processor
In the main text of the paper, we adopt HBT-HERONS (Han et al., 2017; Forouhar Moreno et al., 2025)  [^5] This is an updated version of HBT+ that improves the tracking of subhalos. It can be found at https://github.com/SWIFTSIM/HBT-HERONS as our halo finder, which is a tracking finder. We also explore the result with a phase space finder, VELOCIraptor (Elahi et al., 2019), and compare their results in Appendix A. In this paper, we adopt 20 as the minimum number of particles to be identified as a dark matter halo for both halo finders. We choose the position of the most bound particle as the halo center.
We use SOAP (Spherical Overdensity and Aperture Processor; McGibbon et al. submitted) [^6] https://github.com/SWIFTSIM/SOAP to calculate the spherical overdensity properties for the results from halo finders.
@@ IV.1 Terminology
    where $R$ is the distance to the halo center and $M(<R)$ the enclosed mass that is bound to the halo, within radius $R$. The $R-V_{\rm cir}(R)$ graph is called the galaxy rotation curve, and $V_{\rm max}$ is simply the maximum circular velocity of the rotation curve. In the simulation, the subhalo can lose its mass due to tidal forces from the host halo. However, $V_{\rm max}$ is a more robust measurement of subhalo mass, since it reflects the subhalo’s central, gravitationally bound part, which is less sensitive to how the boundary of the subhalo is defined.
@@ IV.2 Comparison between the PL and the BT models
Figure 4: The comparison between the dark matter mass projection maps with the PL and BT (deep/soft) PPS for m12i at z=0. The color scales are the same across the three panels. All the images are 2D plane projections of cubic volumes with a side length of $400~{}\rm{kpc}$, centered at the main halo’s most bound particle identified by HBT-HERONS.
Table 2 shows some selected properties for the main halos in all the simulations. We learn that the main halo of BT_soft is larger than that of the PL model in mass, radius, concentration, and $V_{\rm max}$. However, the BT_deep model behaves similarly to the PL model, because the mass of the halo is well above the pivot mass of $M_{p}$ based on Equation 6.
@@
In the rest of this subsection, we quantify these differences in terms of the main halo radial density profile, the subhalo mass function, the subhalo $V_{\rm max}$ function, the subhalo radial distribution (subhalo distance function), the cumulative substructure mass fraction and the subhalo $R_{\rm max}-V_{\rm max}$ relationship.
@@ IV.2.1 Main Halo Radial Density Profile
The upper panels of Figure 5 777 For m12f, the main halo center identified by HBT-HERONS is shifted from the real center by $\leq 2{\rm kpc}$, resulting from a bug existing in the code version we used (see Appendix A7 of Forouhar Moreno et al. (2025)), but which has since subsequently been fixed. Given this, we adopt the better centering offered by another halo finder, VELOCIraptor, Elahi et al. (2019) in this figure. Other figures are not affected by this small shift. show a good match between the simulation results and the NFW fitting lines. Thus we can use the concentration parameters given by the NFW profile to describe the radial density profile. The concentration parameters of BT_deep and PL are pretty close (see Table 2), consistent with their similar radial density profiles for the main halos. Meanwhile, the BT_soft model gets a larger concentration parameter than others, consistent with its higher radial density profiles. The unresolved regions are based on the criterion described by subsubsection III.3.1.
@@ IV.2.2 Subhalo Mass Function
Here, $m$ indicates the maximum relative ratio the BT model can reach compared to the PL model, and $\mu_{c}$ is the middle point between $\mu$ starting to drop and $\mu$ value finishing the dropping. According to the definition of Equation 9, $\mu_{c}$ is also where the inverse S shape function drops to half. Since the deep and soft models perform pretty closely in the subhalo mass function, we can use one set of parameters to fit: $m=2,\mu_{c}=10^{-3}$. After getting this fitting function from the bottom ratio panel, we draw it on the upper panel of Figure 6, showing a good fit with the simulation results.
@@ IV.2.4 Subhalo Radial Distribution
As suggested by Springel et al. (2008), the subhalo radial normalized number density in the PL model follows a universal profile across different mass bins, i.e. the grey Aquarius A1 line in the four panels of Figure 8. However, in more massive bins, the subhalos appear to exist only in the outer region. It is likely due to the low number density of subhalos near the center Springel et al. (2008) and lack of statistics using only two zoom-in simulations.
@@ V CONCLUSIONS
JWST’s high redshift observations revealed that galaxies could form earlier than those in the standard model of cosmology Labbé et al. (2023); Boylan-Kolchin (2023); Lovell et al. (2023). The strong lensing observations prefer more substructures Macciò and Miranda (2006); Xu et al. (2009, 2015). Furthermore, the recent observations Müller et al. (2024); Homma et al. (2024) and theories Kim et al. (2018) suggested there could be a “too many satellites” problem in the nearby universe. These motivate us to consider alternative cosmological models with an enhanced Primordial Power Spectrum at small scales ($>1~{}{\rm cMpc}^{-1}$), namely the blue-tilted (BT) model.
In this study, we have conducted cosmological zoom-in simulations of MW host mass halos with the blue-tilted primordial power spectrum. We have considered two blue-tilted models from Ref. Hirano and Yoshida (2024): BT_soft with the enhancement at $\gtrsim 1~{}{\rm cMpc}^{-1}$ and BT_deep with the enhancement at $\gtrsim 4~{}{\rm cMpc}^{-1}$. They are all within the accepted parameter space according to the JWST observation Hirano and Yoshida (2024) and the dwarf galaxy central density observation Dekker and Kravtsov (2024).
@@ Acknowledgments
TKC is supported by the ‘Improvement on Competitiveness in Hiring New Faculties’ Funding Scheme from the Chinese University of Hong Kong (4937210, 4937211, 4937212) and RGC Early Career Scheme (24301524). TKC and JW are partially supported by the Direct Grant project (4053662) from the Chinese University of Hong Kong. VJFM acknowledges support by NWO through the Dark Universe Science Collaboration (OCENW.XL21.XL21.025).
@@ Appendix A Results from another halo finder: VELOCIraptor
Apart from HBT-HERONS, we also identify the halos and subhalos with VELOCIraptor (VR) (Elahi et al., 2019), a code based on the friends-of-friends (FOF) algorithm that identifies peaks in 6D phase-space to find subhalos (both the configuration space and phase space). We compare the HBT-HERONS and VR results on the subhalo mass function, subhalo $V_{\rm max}$ function, and subhalo radial distribution in subsection A.1.
Currently, SOAP (see subsection III.3) only allows the calculation of spherical overdensity properties for field halos. Besides, HBT-HERONS (without SOAP post process) could only calculate limited kinds of SO properties. Hence, we used VR (without SOAP post process) to calculate the virial masses for all the halos (also including subhalos) and show the $M_{\rm vir}-V_{\rm max}$ relationship in subsection A.2. Usually, the spherical overdensity properties will include all of the particles within the SO radius, regardless of whether they are bound to the main halo. However, for subhalos, VR only considers the bound particles [^9] See Mass and radius properties in https://velociraptor-stf.readthedocs.io/en/latest/output.html. Otherwise, the subhalos near the center of the main halo will have unphysically large masses.
Our essential parameters for VR are listed in footnote 12, based on the sample configuration for 6D dark matter-only simulation  [^10] https://velociraptor-stf.readthedocs.io/en/latest/_downloads/97ba16973bcff9d4afe940c2e13d12d2/sample_dmcosmological_6dfof_subhalo.cfg. We set the most bound particle’s position as the reference frame center to be consistent with HBT-HERONS for comparison in subsection A.1.
@@
A.1 Comparison between halo finders: HBT-HERONS and VELOCIraptor
Here, we choose m12i to analyze the difference between the two halo finders. VR stands for VELOCIraptor in both the text and figures. In this subsection, we also use SOAP to process VR and HBT-HERONS catalogs in a consistent manner.
@@ A.1 Comparison between halo finders: HBT-HERONS and VELOCIraptor
Figure 11: The cumulative scaled subhalo mass functions for m12i, with two halo finders HBT-HERONS (colored solid) and VR (colored dashed), respectively. The upper panel shows the subhalo mass function for the PL (orange), BT_deep (green), and BT_soft (red) models, with the subhalo mass $M_{\rm sub}$ scaled by $M_{200c}$ of the main halo. The bottom panel shows the ratios of the VR to HBT-HERONS number. The shaded areas are the same as that in Figure 6. The shaded areas overlap because we set the minimum size of the subhalo to be the same in both halo finders.
The subhalo mass function comparison between VR and HBT-HERONS is shown in Figure 11. These two halo finders agree with each other for $10^{-5}<\mu<10^{-3},10^{7}~{}{\rm M}_{\odot}<M_{\rm sub}<10^{9}~{}{\rm M}_{\odot}$. However, HBT-HERONS is higher and closer to the COCO fitting line than VR at the high mass end, implying that HBT-HERONS is better at finding more massive subhalos. But at the low mass end, VR appears to have more subhalos than HBT-HERONS.
Figure 12: The cumulative scaled subhalo $V_{\rm max}$ functions for m12i, with two halo finders HBT-HERONS (colored solid) and VR (colored dashed), respectively. The upper panel shows the subhalo $V_{\rm max}$ function for the PL (orange), BT_deep (green), and BT_soft (red) models, with the subhalo $V_{\rm max}$ scaled by $V_{200c}$ of the main halo. The bottom panel shows the ratios of the VR to HBT-HERONS number. The shaded areas are the same as those in Figure 7.
The subhalo $V_{\rm max}$ function comparison between VR and HBT-HERONS is shown in Figure 12. These two halo finders give similar results except at the low $V_{\rm max}$ end near the resolution limit. When approaching the limit $V_{\rm max}$ value, VR deviates from the power-law behavior later than HBT-HERONS.
Figure 13: The subhalo radial number density profiles at four different mass bins for m12i, using different halo finders HBT-HERONS (colored solid) and VR (colored dashed). We consider the PL (orange), BT_deep (green), and BT_soft (red) models in each subfigure (also in each mass bin): the upper panel is the same definition as that in Figure 8; the bottom panel shows the ratios of the VR results to the HBT-HERONS results.
The subhalo radial distribution comparison between VR and HBT-HERONS is shown in Figure 13. In mass bins larger than $10^{7}~{}{\rm M}_{\odot}$, VR and HBT-HERONS have nearly the same results. But, at the lower mass end, VR has much fewer subhalos in the inner region than HBT-HERONS and the Aquarius result.
To summarize, VR and HBT-HERONS behave similarly for the subhalo mass and $V_{\rm max}$ functions. However, for the radial properties (the subhalo radial distribution function), VR gaves a lower halo count near the host halo center for the low halo mass bin.